Solve using equation, not a financial calculator. You want to buy a car, and a local bank will lend you $20,000. The loan would be fully

amortized over 5 years (60 months), and the nominal interest rate would be 12%, with

interest paid monthly. What is the monthly loan payment? What is the loan’s EFF%?
Answer #1

Number of periods (n) = 60

Present value of loan (PV) = $20,000

Nominal interest rate = 12% per annum

Monthly interest rate (r) = 12% / 12 = 1% per month.

Using the annuity formula,

PV= PMT* 1-(1+r)^-n/r

PMT = PV*r/1-(1+r)^-n

= 20,000*.01/1-(1+.01)^-60

PMT (The monthly pay) = 444.88

EFF = (1 + APR/12)^12 - 1

= (1 + .12/12)^12 - 1

EFF = 0.1268